The expression for the solubility product of Ba3 (AsO4)2 is?
how and why
Answer #1
We Know that :Ba3(AsO4)2 <---------> 3Ba+2 +2AsO4-3The above salt isA3B2 saltKsp = [3 Ba+2]3 [ 2AsO4-3 ]2= 108S5If we assume the solubility of Ba+2=AsO4-3 = S
answered by: Sylvie